### v1

# Minimal multiplicity of fiber components in abelian fibrations

Abstract. An abelian fibration is a proper projective surjective map of complex varieties with general fiber an abelian variety. Consider a multiple fiber of an abelian fibration, and let $\mu_{1},...,\mu_{k}$ be the multiplicities of its irreducible components. We prove that the minimum of $\mu_{i}$ is equal to their greatest common divisor $\operatorname{\text{\it gcd}}(\mu_{1},...,\mu_{k})$.
1 Introduction
Let $f:X\to B$ be a proper fibration from a complex manifold $X$ onto a smooth complex curve $B$, and $b\in B$ be a point. Let $X_{b}:=f^{*}(b):=\sum_{i}m_{i}F_{i}$ be the scheme theoretic (divisorial, here) fiber of $f$, the positive integers $m_{i}$ being the “multiplicities” of the components $F_{i}$ of $X_{b}$. The “multiplicity” of $X_{b}$ is classically defined as being the greatest common divisor of the numbers $m_{i}$. This is, in particular, the definition used by Kodaira and Shafarevich in their classification of degenerate fibers in elliptic fibrations of surfaces ([BHPV]). For higher dimensional bases $B$, one looks at the generic fibers over the components of the discriminant divisor of $f$ in $B$.
In [Ca2], however, a different definition is used: one does not consider there the $\operatorname{\text{\it gcd}}$ of the $m_{i}$, but their infimum. This second definition is dictated by the need to characterize what is called there the “Bogomolov sheaves” on projective or compact Kähler varieties $X$, which are the rank-one subsheaves $L\subset\Omega^{p}_{X},p>0$ such that $\kappa(X,L)=p$, the top Kodaira dimension, after Bogomolov’s bound ([Bo, CCR]). These sheaves, analogs of “big” canonical bundles, indeed play a central rôle in bimeromorphic classification (see [Ca2] for more details).
It is thus important to know when the two definitions agree (for locally projective or Kähler proper fibrations). This is the case when the smooth fibers of $f$ are either rationally connected (by [GHS] or [JS, main theorem]), or are compact complex tori (in particular: Abelian varieties), which is the topic of the present text. In particular, the two definitions agree for the elliptic surfaces of Kodaira/Shafarevich’s classification.
On the other hand, examples of fibrations with fibers higher genus curves for which the two versions differ are found in [W], [S], [Ca3]. A remarkable example is given in [L], where the smooth fibers are two Enriques surfaces of multiplicities 2 and 3, with consequences in hyperbolic and arithmetic geometries derived in [BCJW].
In the Appendix 4 we introduce the abelian torsor structure in the settings of complex torus fibrations, as in [BKV, KV]. We use the averaging formalism applied to multisections of an abelian fibration in order to obtain the following theorem over a $1$-dimensional base.
Theorem 1.1: (2) Let $\pi:\;M\to B$ be an abelian fibration, that is, a proper projective surjective map of complex varieties with general fiber an abelian variety. Assume that all points of $B$ are regular values of $\pi$ except $x_{0}\in B$, and $\dim_{\mathbb{C}}B=1$. Let $F_{1},...,F_{k}$ be the irreducible components of $\pi^{-1}(x_{0})$, and $\mu_{1},...,\mu_{k}$ their multiplicities. Assume that the greatest common divisor of $\mu_{i}$ is 1. Then $\pi$ admits a local section.
In Section 3, we generalize these results to higher-dimensional base.
Theorem 1.2: (3) Let $f:M\to B$ be a proper Kähler fibration with $M$ smooth, $B$ normal, such that its generic fibers are smooth complex tori. Assume that $f$ has equidimensional fibers. Then for any $b\in B$, $f$ has a meromorphic section above some open neighborhood of $b$ in $B$.
Corollary 1.3: (3) In the situation of 3, assume that for some $b\in B$, the fiber is (in the sense of [BM]) $M_{b}=\sum_{i}m_{i}F_{i}$, with $F_{i}$ being the irreducible components. If $gcd_{i}(\{m_{i}\})=1$, then $m_{i}=1$ for some $i$.
Remark 1.4: In assumptions of 3, it is essential that the fibration is (at least locally) Kähler: one of the key arguments in Section 3 uses the compactness of the Barlet spaces. However, in the settings of hyperkähler geometry, the first-named author proved that Lagrangian fibrations on hyperkähler manifolds are not only Kähler, but also locally projective ([Ca1]).
2 Averaging multisections in a torus fibration over a disk
In Section 4, we define the torsors and described the averaging formalism ([BKV]), which is a general method of constructing a section of a torsor bundle, starting with a (signed) multisection of total multiplicity 1. Applying this construction to an abelian fibration, we obtain the following theorem.
Theorem 2.1: Let $\pi:\;M\to B$ be a torus fibration, that is, a proper map of complex varieties with general fiber a compact torus. Assume that all points of $B$ are regular values of $\pi$ except $x_{0}\in B$, and $\dim_{\mathbb{C}}B=1$. Let $F_{1},...,F_{k}$ be the irreducible components of $\pi^{-1}(x_{0})$, and $\mu_{1},...,\mu_{k}$ their multiplicities. Assume that the greatest common divisor of $\mu_{i}$ is 1. Then $\pi$ admits a local section, and ${\it min}_{i}\mu_{i}=1$.
Proof. Step 1: Around a smooth point $m\in F_{i}$, the projection maps can be written in local coordinates as $(z_{1},z_{2},...,z_{n})\mapsto z_{1}^{\mu_{i}}$. Let ${\cal E}_{\mu_{i}}$ be the set of all roots of unity of degree $\mu_{i}$. Then the image of the multivalued map $t\to({\cal E}_{\mu_{i}}t^{\mu_{i}^{-1}},0,...,0)$ is a smooth curve in $M$, and its projection to a neighbourhood $U$ of $x\in B$ is a $\mu_{i}$-sheeted ramified cover.
Step 2: Consider such a multisection $Z_{i}$ for each $F_{i}$, and let $a_{1},...,a_{k}$ be a set of integers such that $\sum_{i}a_{i}\mu_{i}=1$. We consider $\sum a_{i}Z_{i}$ as a multisection with multiplicites determined by $a_{i}$, with total multiplicity 1. Applying the averaging construction (4.2), we obtain a meromorphic map $\varphi:\;U\backslash x_{0}\to M$, which is holomorphic and Lipschitz by construction. By Riemann removable singularity theorem, this map extends to the origin $x_{0}\in U$.
Corollary 2.2: Let $\pi:\;M\to B$ be an torus fibration, that is, a proper surjective map of complex varieties with general fiber a compact torus. Assume that all points of $B$ are regular values of $\pi$ except $x_{0}\in B$, and $\dim_{\mathbb{C}}B=1$. Let $F_{1},...,F_{k}$ be the irreducible components of $\pi^{-1}(x_{0})$, and $\mu_{1},...,\mu_{k}$ their multiplicities. Then ${\it min}\mu_{i}=\operatorname{\text{\it gcd}}(\mu_{i})$, where $\operatorname{\text{\it gcd}}$ denotes the greatest common divisor.
Proof: Passing to a ramified covering of $X$ of order $\operatorname{\text{\it gcd}}(\mu_{i})$, we arrive to an torus fibration with the same fibers with multiplicities $\mu_{i}^{\prime}=\frac{\mu_{i}}{\operatorname{\text{\it gcd}}(\mu_{i})}$, hence $\operatorname{\text{\it gcd}}(\mu_{i}^{\prime})=1$. By 2, it has a section. This section gives a multisection of order $\operatorname{\text{\it gcd}}(\mu_{i})$ of $\mu$.
3 Higher dimensional bases
The aim of this section is to prove the following results generalizing 2 from a $1$-dimensional base to a base of any dimension. First we give a proof of the corollary, and then of the theorem below.
Theorem 3.1: Let $f:M\to B$ be a proper Kähler fibration between normal complex spaces such that its generic fibers are smooth compact complex tori. Assume that $f$ has equidimensional fibers. Then for any $b\in B$, $f$ has a meromorphic section above some open neighborhood of $b$ in $B$.
Corollary 3.2: In the situation of 3, assume that for some $b\in B$, the fiber is (in the sense of [BM]) $M_{b}=\sum_{i}m_{i}F_{i}$, with $F_{i}$ being the irreducible components. If $gcd_{i}(\{m_{i}\})=1$, then ${\it min}_{i}(\{m_{i}\})=1$.
Proof of 3, assuming 3: Let $s:B\dasharrow M$ be a meromorphic section of $f$ defined near $b$. Let $B^{\prime}$ be the closure of the image of the meromorphic section $B\dasharrow M$. Let $m:B^{\prime}\to B$ be the proper bimeromorphic map such that $s^{\prime}:=s\circ m:B^{\prime}\to M^{\prime}:=M\times_{B}B^{\prime}$ is a holomorphic section of $f^{\prime}:M^{\prime}\to B^{\prime}$ obtained from $f$ by the base-change $m:B^{\prime}\to B$. For any $b^{\prime}\in B^{\prime}$ such that $m(b^{\prime})=b$, the component of $M^{\prime}_{b^{\prime}}$ containing $s^{\prime}(b^{\prime})$ has multiplicity $1$ in $M^{\prime}_{b^{\prime}}=M_{b}\times_{\{b\}}\{b^{\prime}\}$. Thus, so has the corresponding component of $M_{b}$: since the existence of a local holomorphic section implies that the multiplicity of the corresponding component of the fiber has multiplicity one.
Proof of 3: Let $B^{0}\subset B$ be the Zariski open subset over which the fibers are smooth compact tori. Fix $b\in B^{0}$, let $A_{b}=\operatorname{Aut}^{0}(M_{b})$ be the translation group of the fiber $M_{b}$. Clearly, $A_{b}$ is a torus, isomorphic to $M_{b}$. If $u_{i}\in\mathbb{Z},i=1,\dots,r$ is such that $\sum_{i}u_{i}=1$ and if $p_{i}\in M_{b},i=1,\dots,r$, the sum $\sum_{i}u_{i}\cdot p_{i}=s\in M_{b}$ is well-defined, independently of the choice of an origin in $M_{b}$ (4.2).
Recall ([BM], chap IV, Théorème 3.4.1, Corollaire 3.4.2) that $f$ being equidimensional, with $B$ normal, the components of the fibers of $f$ can be given in a unique way multiplicities (equal to $1$ for those above $B^{0}$) in such a way that the fibers of $f$ form an analytic family of cycles in the Barlet sense, such that the map $\beta:B\to C(M/B)$ sending $M_{b}$ to its reprensentative point in $C(M/B)$, the Chow/Barlet space of relative $f$-cycles of $M$, is analytic. Moreover, if $C:=\beta(B)\subset C(M/B)$, and if $\Gamma_{C}\subset C\times M$ is the incidence graph of the universal family of cycles parametrised by $C$, then $(f\times id_{X})(M)=(\beta\times id_{X})^{-1}(\Gamma_{C})$: by its very definition, the graph of $f$ is obtained from $\Gamma_{C}$ by the base-change $\beta$.
Let now $b^{\prime}\in B\setminus B^{0}$, and $M_{b^{\prime}}=\sum_{i}m_{i}\cdot F_{i}$ be its $f$-fiber. By normality of $M$, through the generic point of each $F_{i}$ we have a local multisection $s_{i},i=1,\dots,r$ of $f$ which has degree $m_{i}$ over $B$. Denote with $p_{i,j}(b)$ its points over $b$. By hypothesis, the gcd of the numbers $\{m_{i}\}$ is one, and Bezout lemma implies the existence of integers $v_{i}$ such that $\sum_{i}u_{i}:=\sum_{i}v_{i}\cdot m_{i}=1,u_{i}:=v_{i}\cdot m_{i}$. By averaging (4.2), for each $b\in B^{0}$ near $b^{\prime}$, a well-defined point $s(b):=\sum_{i,j}u_{i}\cdot p_{i,j}(b)\in M_{b}$, which depends holomorphically on $b$.
We shall show, using the Kählerianity assumption, that this section of $f$ over $B^{0}$ near $b^{\prime}$ extends meromorphically to $B$ near $b^{\prime}$. More precisely, we will show that the addition (and subtraction) in the generic smooth fibers of $f$ extend meromorphically to $M\times_{B}M$. Any section obtained by averaging over the locus $B^{0}$ of smooth fibers of $f$ in $B$ thus extends meromorphically to $B$.
This extension of addition/substraction is based on the properness of the irreducible components of the relative Chow/Barlet space $C(M/B)$ (see [BM] and the references there). Let $B^{0}\subset B$ be the Zariski open subset over which the fibers are smooth compact tori, and let $M^{0}:=f^{-1}(B^{0})$.
Over $B^{0}$ one can define the unit component $\operatorname{Aut}^{0}(M^{0}/B^{0})\to B^{0}$ of the group of translations in the fibers of $f$ as the component of $C(M^{0}\times_{B^{0}}M^{0})\to B^{0}$ consisting of the graphs of these translations. Over $B^{0}$, the incidence graph $\Phi^{0}\subset(\operatorname{Aut}^{0}(M^{0}/B^{0})\times_{B^{0}}(M^{0}\times_{B^{0}}M^{0})$ of the family of relative translations in the fibers of $f$ is also nothing but the graph of the addition map $\operatorname{Aut}^{0}(M^{0}/B^{0})\times M^{0}\to M^{0}$ over $B^{0}$ sending $(t_{b},x_{b})\to t_{b}+x_{b}$, the substraction map being defined as exchanging the $M^{0}/B^{0}$ factors. Since over $B^{0}$ these two maps just consist in exchanging the suitable factors, which obviously extends meromorphically over $B$, they extend meromorphically over $B$ to maps: $\operatorname{Aut}^{0}(M/B)\times_{B}M_{b}$, where $Aut^{0}(M/B)$ is the Zariski closure in $C(M/B)$ of the unit component $\operatorname{Aut}^{0}(M^{0}/B^{0})$. If $u_{i}\in\mathbb{Z},i=1,\dots r$ are such that $\sum_{i}u_{i}=1$, the well-defined map over $B^{0}$: $M^{0}\times_{B^{0}}\times\dots_{B^{0}}M^{0}\to M_{0}$ (r-factors) sending $(x_{1,b},\dots x_{r,b})\to\sum_{i}u_{i}\cdot x_{i,b}$ thus extends meromorphically to $M_{B}\times\dots_{B}M$ over $B$, which shows the claim.
Remark 3.3: 1. Constructions similar to the one given by Ulf Persson in [P, p.83], show that the relative Kählerianity of $f$ is essential for the existence of this meromorphic section.
2. The hypothesis on the $\operatorname{\text{\it gcd}}$ of the $m_{i}$ being equal to $1$ do not follow from the same property in codimension $1$, even if $M$ and $B$ are smooth, $f$ projective, and if the fibers outside $b\in B$ are reduced, as shown by the example of I. Hellmann ([Hel, Section 2.2]).
4 Appendix: Torsor fibrations
4.1 Torsor structure in complex torus fibrations
By definition, a torsor over a group $G$ is a set where $G$ acts freely and transitively. However, the torsor structure can be defined without referring to an a priori given group structure as follows.
Following [KV], we define an abelian torsor structure as a kind of algebraic structure, without first defining the group acting on this torsor. This simplifies some of the definitions when the torsor structure is defined in families, and we do not have the corresponding family of abelian groups on hand.
In the sequel we will make a distinction between the compact torus $T$ and the same torus equipped with a group structure, which we denote $\operatorname{Alb}(T)$. This group structure is clearly unique up to a parallel transport, that is, determined by the choice of 0. The torus $\operatorname{Alb}(T)$ is identified with the abelian variety of $T$.
Definition 4.1: An abelian torsor is a set $S$ equipped with a map $\mu:\;S\times S\times S\to S$ such that
for any $e\in S$, the map $x\cdot y:=\mu(x,y,e)$ defines a commutative and associative product on $S$ with $e$ an identity element
for any fixed $y\in S$, the map $x\mapsto x\cdot y$ is bijective.
Remark 4.2: From the axioms of an abelian torsor, it is clear that the operation $\cdot$” defines a structure of an abelian group on $S$.
Remark 4.3: The identity element axiom means that $\mu(x,y,y)=x$ for any $x$ and $y$. Given a group structure on $S$, the torsor operation $\mu$ can be defined as $\mu(x,y,z)=x+y-z$.
Claim 4.4: Let $\pi:\;M\to B$ be a surjective holomorphic fibration whose general fiber is a compact complex torus. Then the fibers $F_{x}:=\pi^{-1}(x)$ of $\pi$ are equipped with a holomorphic torsor structure which depends on $x\in B$ holomorphically.
Proof: It is not hard to see that any holomorphic map of a compact complex torus is affine, that is, obtained as a composition of a parallel transport and a group homomorphism. Therefore, the holomorphic torsor structure on a compact torus $T$ is unique; indeed, the connected component of the group of automorphisms of $T$ is $\operatorname{Alb}(T)$. Taking a local section, we can identify each fiber $F_{x}$ with its Albanese manifold $\operatorname{Alb}(F_{x})$ holomorphically, and this turns $F_{x}$ into an $\operatorname{Alb}(F_{x})$-torsor. The dependence on $x$ is holomorphic by construction, and it is independent from the choice of a local section by uniqueness.
4.2 The averaging formalism
An affine space is, by definition, a torsor over a vector space. Consider an affine space $A$ over a vector space $V$. After choosing the origin $a\in A$, we can identify $A$ and $V$. For any collection of points $x_{1},...,x_{n}\in A$, their mass center $a+\frac{1}{n}\sum_{i}(x_{i}-a)$ is well defined and independent from the choice of the origin.
This construction works for torsors over infinitely divisible, torsion-free abelian groups (that is, vector spaces over $\mathbb{Q}$}). This assumption is essential, because if the group has torsion, the mass center becomes ambiguous.
However, a version of this construction survives. It was used in [BKV] to produce a rational section of a Lagrangian fibration in a hyperkähler manifold. If we take a weighted sum of $m$ points with positive sign and (m1)(m-1) points with negative sign, the sum is independent from the choice of an origin. We give an abstract version of this construction here, see [BKV, Claim 3.6] for its proof.
Claim 4.5: Let $X$ be a torsor over $G$. Choosing an origin $e\in X$, we can identify $X$ with $G$. Then, for any $m\in{\mathbb{Z}}^{>0}$ there is a natural map $X^{m}\times X^{m-1}\to X$ taking $(x_{1},...,x_{m},y_{1},...,y_{m-1})$ to $e+\sum_{i=1}^{m}(x_{i}-e)-\sum_{i=1}^{m-1}(y_{i}-e)$, which is independent from the choice of $e$.
Acknowledgments. L. K. and M. V. are grateful to J. Kollár for his insightful comments and email exchanges, and to F. Bogomolov, whose ideas were of great influence in this paper.
The present text was written during a visit of F. C. at IMPA in Rio, and he thanks this institute and the local branch of CNRS for their support, making this visit possible.